OpenStudy (narissa):

the ratio of the perimeters of 2 rectangles is 4 to 7 the perimeter of the larger rectangle is 42 what is the ratio for the smaller one?

OpenStudy (narissa):

would it be 24?

OpenStudy (anonymous):

i think soo. yes

OpenStudy (anonymous):

then the smaller one would be 8 to 16

OpenStudy (narissa):

I muliply 6*7 and 4*6 is that how u do it?

OpenStudy (anonymous):

idk. but let me do some research then ill tell you step by step for the equation.

OpenStudy (narissa):

k

OpenStudy (anonymous):

The ratio of the perimeters is equal to the scale factor. If rectangle #1 has sides L and W, then the perimeter is 2*L1 + 2*W1 = 2*(L1 + W1). If rectangle # 2 is similar to #1 and sides are scaled by a factor S, so that L2 = S*L1 and W2 = S*W1, the perimeter of rectangle #2 is 2*(L2 + W2)= 2*(S*L1 + S*W1) = S*2*(L1 + W1) = S*(perimeter of rectangle #1)

OpenStudy (anonymous):

my research has lead me to this information. but idk if it helps you. :(

OpenStudy (narissa):

thats confusing!!!!

OpenStudy (anonymous):

yep it is??..?? but i think you are right for your answer. idk youll have to ask math experts on this. im just a intermediate at math.

OpenStudy (narissa):

well thanks anyway

OpenStudy (anonymous):

24 is right answer.